# SpecDetect: Simple, Fast, and Training-Free Detection of LLM-Generated Text via Spectral Analysis

### v2

@@ Abstract
The proliferation of high-quality text from Large Language Models (LLMs) demands reliable and efficient detection methods. While existing training-free approaches show promise, they often rely on surface-level statistics and overlook fundamental signal properties of the text generation process. In this work, we reframe detection as a signal processing problem, introducing a novel paradigm that analyzes the sequence of token log-probabilities in the frequency domain. By systematically analyzing the signal’s spectral properties using the global Discrete Fourier Transform (DFT) and the local Short-Time Fourier Transform (STFT), we find that human-written text consistently exhibits significantly higher spectral energy. This higher energy reflects the larger-amplitude fluctuations inherent in human writing compared to the suppressed dynamics of LLM-generated text. Based on this key insight, we construct SpecDetect, a detector built on a single, robust feature from the global DFT: DFT total energy. We also propose an enhanced version, SpecDetect++, which incorporates a sampling discrepancy mechanism to further boost robustness. Extensive experiments demonstrate that our approach outperforms the state-of-the-art model while running in nearly half the time. Our work introduces a new, efficient, and interpretable pathway for LLM-generated text detection, showing that classical signal processing techniques offer a surprisingly powerful solution to this modern challenge.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have enabled the generation of high-quality text that is often indistinguishable from human writing (Crothers, Japkowicz, and Viktor 2023). While beneficial, this capability poses significant challenges related to potential misuse, including the spread of misinformation (Opdahl et al. 2023; Fang et al. 2024) and threats to academic integrity (Else 2023; Currie 2023), which underscore the urgent need for reliable LLM-generated text detection methods.
@@ 3.1 Frequency-Domain Analysis
To systematically explore spectral differences between human and LLM-generated text, we analyze three datasets: XSum (Narayan, Cohen, and Lapata 2018), Reddit ELI5 (Fan et al. 2019), and WritingPrompts (Fan, Lewis, and Dauphin 2018). Using source models GPT-J (Wang et al. 2023), LLaMA3-8B (AI@Meta 2024), and GPT-4-Turbo (Achiam et al. 2024) (with GPT-J as proxy), we test both white-box (GPT-J) and black-box (LLaMA3-8B, GPT-4-Turbo) conditions. Figure 1 shows a sample pair for intuition (STFT visualizations in Appendix Figure 10), suggesting human text has a more volatile time-domain signal, with a spectrum of significantly higher overall energy.
@@ 4.1 Experimental Setup
We evaluate our methods on a diverse suite of datasets, following the configurations of prior work (Bao et al. 2023; Xu et al. 2024). Our experiments are conducted in two settings: a white-box scenario using the XSum (Narayan, Cohen, and Lapata 2018), SQuAD (Rajpurkar et al. 2016), and WritingPrompts (Fan, Lewis, and Dauphin 2018) datasets, and a more realistic black-box scenario using XSum, Reddit ELI5 (Fan et al. 2019), and WritingPrompts. To assess performance on other languages, our analysis also includes two non-English datasets: WMT16-De (German) (Bojar et al. 2016) and Zhihu-Economy (Chinese) (Xu et al. 2024). In all cases, LLM-generated text is created by prompting source models with the initial 30 tokens of each human-written example, ensuring the human and machine-generated texts have matching lengths. More details are in Appendix A.2.
@@
We compare our approach against 10 training-free baselines, which are divided into two groups (see Appendix A.1 for full details). Sample-based methods directly use token statistics for scoring and include Log-Likelihood (Solaiman et al. 2019), LogRank (Solaiman et al. 2019), Entropy (Gehrmann, Strobelt, and Rush 2019; Ippolito et al. 2019), DetectLRR (Su et al. 2023), and Lastde (Xu et al. 2024). In contrast, distribution-based methods score text by contrasting it with generated samples and include DetectGPT (Mitchell et al. 2023), DetectNPR (Su et al. 2023), DNA-GPT (Yang et al. 2023), Fast-DetectGPT (Bao et al. 2023), and Lastde++ (Xu et al. 2024). Consistent with prior work (Xu et al. 2024), we use AUC as the evaluation metric and set the sampling number for SpecDetect++ to 100.
@@ 4.3 Robustness and In-depth Analysis
We evaluate our method’s robustness across challenging scenarios. To reflect real-world conditions, all analyses in this section are conducted in a black-box setting, focusing on the comparison with high-performing baselines. Beyond the main experiments, Appendix B.7 and B.8 include more analyses: impacts of decoding strategies and performance of alternative energy-based features as detectors.
@@
The choice of proxy model is a critical factor influencing black-box detection performance. We investigate this by evaluating detectors on the XSum dataset across four different source-proxy pairs, with results presented in Figure 8.
@@
The results confirm that performance varies significantly with the proxy model, highlighting the difficulty of black-box detection. However, across all tested pairs, our methods demonstrate superior generalization. Our base method consistently and significantly outperforms baselines by a large margin. This indicates that our fundamental spectral energy feature is more robust to proxy model mismatch than complex time-series features. Similarly, our enhanced method SpecDetect++ consistently outperforms other distribution-based methods, showcasing its robustness.
@@
Figure 9: Detection performance (AUC) in two non-English scenarios: a German task on the WMT16-De dataset (left y-axis; source: mGPT, proxy: GPT-J) and a Chinese task on the Zhihu-Economy dataset (right y-axis; source: Qwen-8B, proxy: Yi-1.5-6B).
To evaluate our method’s generalizability beyond English, we conduct experiments on German and Chinese datasets, with results shown in Figure 9. Following previous work (Xu et al. 2024), we test two scenarios: a German task on the WMT16 dataset using mGPT as the source model and GPT-J as the proxy model, and a Chinese task on an economics dataset using Qwen-7B as the source and Yi-1.5-6B as the proxy. The results show the strong performance of our framework across these different languages. Our base method, SpecDetect, proves to be highly effective, achieving a top-tier AUC. Moreover, our enhanced version, SpecDetect++, also delivers very competitive results, outperforming the state-of-the-art Lastde++ in the German scenario. This highlights the effectiveness of our approach in non-English settings.
@@ 5 Conclusion
In this work, we introduce SpecDetect, a novel, training-free detector for LLM-generated text that operates from a frequency-domain perspective. By treating the token probability sequence as a signal, we demonstrate that the fundamental difference between human and machine text generation can be robustly captured by a single, hyperparameter-free feature: the DFT Total Energy. Our proposed method is conceptually simple and computationally efficient. Extensive experiments confirm superiority: SpecDetect and SpecDetect++ set a new state-of-the-art in effectiveness-efficiency trade-off, with superior or comparable performance to the latest models at half computational cost and strong robustness. This work advances LLM text detection via classical signal processing for NLP challenges. Future research may explore more spectral features for fine-grained tasks like model attribution.
@@ A.1 Baseline Descriptions
  • Entropy (Gehrmann, Strobelt, and Rush 2019; Ippolito et al. 2019): The metric is the average token entropy, calculated from the probability distribution over the entire vocabulary at each position.
@@ A.2 Dataset Details
Following the configurations in prior studies (Bao et al. 2023; Yang et al. 2023; Xu et al. 2024), we evaluate our approach using four English datasets: (i) XSum (Narayan, Cohen, and Lapata 2018), which contains BBC News articles for summarization tasks; (ii) SQuAD (Rajpurkar et al. 2016; Rajpurkar, Jia, and Liang 2018), comprising Wikipedia-based question-answering contexts; (iii) WritingPrompts (Fan, Lewis, and Dauphin 2018), featuring creative story generation prompts; and (iv) Reddit ELI5 (Fan et al. 2019), an explanatory question-answering dataset covering biology, physics, chemistry, economics, law, and technology. To assess cross-lingual robustness, following (Xu et al. 2024), we include two non-English datasets: (i) WMT16-De (Bojar et al. 2016), a German dataset originally used for translation tasks; and Zhihu-Economy(Xu et al. 2024), a Chinese question-answering dataset focusing on economics. Following previous work, we use Xsum, SQuAD, and WritingPrompt for white-box detection, while Xsum, Reddit ELI5, and WritingPrompts for black-box detection. Each dataset includes 150 human-written examples. For LLM-generated text, we use the first 30 tokens of each human-written example as a prompt, with continuations generated by source models. As a result, each finalized dataset comprises 150 negative samples (human-written) and 150 positive samples (LLM-generated), with each pair featuring the same text length and matching initial segments.
@@ B.6 Impact of Contrast Sample Size
We evaluate the impact of different decoding strategies on detection performance. Figure 11 presents the average AUC scores of various detectors across texts generated by three source models (GPT-2, Neo-2.7B, OPT-2.7B) using top-k (Fan, Lewis, and Dauphin 2018), top-p (Holtzman et al. 2019), and temperature sampling (Ackley, Hinton, and Sejnowski 1985). Specifically, following the setup in previous work (Xu et al. 2024), we configure the parameters as: top-p = 0.96, top-k = 40, and temperature = 0.80.

### v3

@@ Abstract
The proliferation of high-quality text from Large Language Models (LLMs) demands reliable and efficient detection methods. While existing training-free approaches show promise, they often rely on surface-level statistics and overlook fundamental signal properties of the text generation process. In this work, we reframe detection as a signal processing problem, introducing a novel paradigm that analyzes the sequence of token log-probabilities in the frequency domain. By systematically analyzing the signal’s spectral properties using the global Discrete Fourier Transform (DFT) and the local Short-Time Fourier Transform (STFT), we find that human-written text consistently exhibits significantly higher spectral energy. This higher energy reflects the larger-amplitude fluctuations inherent in human writing compared to the suppressed dynamics of LLM-generated text. Based on this key insight, we construct SpecDetect, a detector built on a single, robust feature from the global DFT: DFT total energy. We also propose an enhanced version, SpecDetect++, which incorporates a sampling discrepancy mechanism to further boost robustness. Extensive experiments show that our approach outperforms the state-of-the-art model while running in nearly half the time. Our work introduces a new, efficient, and interpretable pathway for LLM-generated text detection, showing that classical signal processing techniques offer a surprisingly powerful solution to this modern challenge.
Codehttps://github.com/luohaitong/SpecDetect
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have enabled the generation of high-quality text that is often indistinguishable from human writing (Crothers et al. 2023). While beneficial, this capability poses significant challenges related to potential misuse, including the spread of misinformation (Opdahl et al. 2023; Fang et al. 2024) and threats to academic integrity (Else 2023; Currie 2023), which underscore the urgent need for reliable LLM-generated text detection methods.
@@ 3.1 Frequency-Domain Analysis
To systematically explore spectral differences between human and LLM-generated text, we analyze three datasets: XSum (Narayan et al. 2018), Reddit ELI5 (Fan et al. 2019), and WritingPrompts (Fan et al. 2018). Using source models GPT-J (Wang et al. 2023), LLaMA3-8B (AI@Meta 2024), and GPT-4-Turbo (Achiam et al. 2024) (with GPT-J as proxy), we test both white-box (GPT-J) and black-box (LLaMA3-8B, GPT-4-Turbo) conditions. Figure 1 shows a sample pair for intuition (STFT visualizations in Appendix Figure 10), suggesting human text has a more volatile time-domain signal, with a spectrum of significantly higher overall energy.
@@ 4.1 Experimental Setup
We evaluate our methods on a diverse suite of datasets, following the configurations of prior work (Bao et al. 2023; Xu et al. 2024). Our experiments are conducted in two settings: a white-box scenario using the XSum (Narayan et al. 2018), SQuAD (Rajpurkar et al. 2016), and WritingPrompts (Fan et al. 2018) datasets, and a more realistic black-box scenario using XSum, Reddit ELI5 (Fan et al. 2019), and WritingPrompts. To assess performance on other languages, our analysis also includes two non-English datasets: WMT16-De (German) (Bojar et al. 2016) and Zhihu-Economy (Chinese) (Xu et al. 2024). In all cases, LLM-generated text is created by prompting source models with the initial 30 tokens of each human-written example, ensuring the human and machine-generated texts have matching lengths. More details are in Appendix A.2.
@@
We compare our approach against 10 training-free baselines, which are divided into two groups (see Appendix A.1 for full details). Sample-based methods directly use token statistics for scoring and include Log-Likelihood (Solaiman et al. 2019), LogRank (Solaiman et al. 2019), Entropy (Gehrmann et al. 2019; Ippolito et al. 2019), DetectLRR (Su et al. 2023), and Lastde (Xu et al. 2024). In contrast, distribution-based methods score text by contrasting it with generated samples and include DetectGPT (Mitchell et al. 2023), DetectNPR (Su et al. 2023), DNA-GPT (Yang et al. 2023), Fast-DetectGPT (Bao et al. 2023), and Lastde++ (Xu et al. 2024). Consistent with prior work (Xu et al. 2024), we use AUC as the evaluation metric and set the sampling number for SpecDetect++ to 100.
@@ 4.3 Robustness and In-depth Analysis
We evaluate our method’s robustness in challenging, black-box scenarios. Further analyses on decoding strategies and alternative detectors are in Appendix B.7 and B.8.
@@
Proxy model choice critically impacts black-box detection. We investigate this by evaluating detectors on XSum across four different source-proxy pairs, with results in Figure 8. While performance varies significantly, confirming the challenge, our methods demonstrate superior generalization. Our SpecDetect consistently outperforms baselines by a large margin, indicating its spectral feature is more robust to proxy mismatch than complex time-series features. Similarly, SpecDetect++ consistently outperforms other distribution-based methods, showcasing its robustness across all tested pairs.
@@
Figure 9: Performance in two non-English scenarios: a German task on WMT16-De (left y-axis; source: mGPT, proxy: GPT-J) and a Chinese task on Zhihu-Economy (right y-axis; source: Qwen-8B, proxy: Yi-1.5-6B).
To evaluate generalizability beyond English, we conduct German and Chinese experiments, shown in Figure 9. Following previous work (Xu et al. 2024), we test German on WMT16 (source: mGPT, proxy: GPT-J) and Chinese on an economics dataset (source: Qwen-7B, proxy: Yi-1.5-6B). The results show our framework’s strong cross-lingual performance. Our base method, SpecDetect, achieves top-tier AUC. The enhanced SpecDetect++ is also highly competitive, outperforming the SOTA Lastde++ in the German scenario, highlighting our approach’s effectiveness in non-English settings.
@@ 5 Conclusion
We introduce SpecDetect, a novel, training-free LLM text detector using a frequency-domain perspective. By treating the token probability sequence as a signal, it uses a single, hyperparameter-free feature, DFT Total Energy, to robustly distinguish human from machine text. This method is simple and efficient. Experiments confirm SpecDetect set a new SOTA in effectiveness-efficiency trade-off, achieving superior or comparable performance at half the computational cost with strong robustness. This work advances LLM text detection via classical signal processing, and future research can explore more spectral features for fine-grained tasks.
Acknowledgments
This work was supported in whole or in part, by National Natural Science Foundation of China (U24B6012, U2333201, 62372429), the Innovation Funding of ICT, CAS under Grant No. E461040, Pilot for Major Scientific Research Facility of Jiangsu Province of China (No.BM2021800).
@@ A.1 Baseline Descriptions
  • Entropy (Gehrmann et al. 2019; Ippolito et al. 2019): The metric is the average token entropy, calculated from the probability distribution over the entire vocabulary at each position.
@@ A.2 Dataset Details
Following the configurations in prior studies (Bao et al. 2023; Yang et al. 2023; Xu et al. 2024), we evaluate our approach using four English datasets: (i) XSum (Narayan et al. 2018), which contains BBC News articles for summarization tasks; (ii) SQuAD (Rajpurkar et al. 2016, 2018), comprising Wikipedia-based question-answering contexts; (iii) WritingPrompts (Fan et al. 2018), featuring creative story generation prompts; and (iv) Reddit ELI5 (Fan et al. 2019), an explanatory question-answering dataset covering biology, physics, chemistry, economics, law, and technology. To assess cross-lingual robustness, following (Xu et al. 2024), we include two non-English datasets: (i) WMT16-De (Bojar et al. 2016), a German dataset originally used for translation tasks; and Zhihu-Economy(Xu et al. 2024), a Chinese question-answering dataset focusing on economics. Following previous work, we use Xsum, SQuAD, and WritingPrompt for white-box detection, while Xsum, Reddit ELI5, and WritingPrompts for black-box detection. Each dataset includes 150 human-written examples. For LLM-generated text, we use the first 30 tokens of each human-written example as a prompt, with continuations generated by source models. As a result, each finalized dataset comprises 150 negative samples (human-written) and 150 positive samples (LLM-generated), with each pair featuring the same text length and matching initial segments.
@@ B.6 Impact of Contrast Sample Size
We evaluate the impact of different decoding strategies on detection performance. Figure 11 presents the average AUC scores of various detectors across texts generated by three source models (GPT-2, Neo-2.7B, OPT-2.7B) using top-k (Fan et al. 2018), top-p (Holtzman et al. 2019), and temperature sampling (Ackley et al. 1985). Specifically, following the setup in previous work (Xu et al. 2024), we configure the parameters as: top-p = 0.96, top-k = 40, and temperature = 0.80.